40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8870     Entry time: Thu Jul 18 15:34:15 2013
Author: Alex Cole 
Type: Update 
Category: Electronics 
Subject: PD Frequency Response Update 

 [Eric, Alex]

Our RF Switch arrived today, and we mounted it in rack 1Y1 (1st attachment). 

We connect our input fiber and all of our output fibers to our 1x16 optical splitter (2nd attachment). Note that the 75 meter fiber we are using for the splitter's input is in a very temporary position (3rd attachment - it's the spool).

We successfully turned our laser on and tested the optical splitter by measuring output power at each fiber using our Thorlabs PM20 power meter. Data was taken with the laser running at 67.5 mA and 24 degrees Celsius:

Detector name                  Power

REF DET 192 µW
AS55 146 µW
REFL55 180 µW
REFL11 172 µW
MCREFL 133 µW
REFL33 146 µW
REFL165 180 µW
POP22/POP110 182 µW
POP55 193 µW
POX11 123 µW

 

 

Attachment 1: photo_3_(1).JPG  137 kB  | Show | Show all
Attachment 2: photo_1_(4).JPG  126 kB  | Show | Show all
Attachment 3: photo_2_(4).JPG  118 kB  | Show | Show all
ELOG V3.1.3-